Cleveland Cove Enterprises is evaluating the purchase of an elaborate hydraulic lift system for all of its locations to use for the boats brought in for repair. The company has narrowed their choices down to​ two: the B14 Model and the F54 Model. Financial data about the two choices follows. ​(For the B14​ Model, the present value in this scenario is

0.507

and the present value of annuity in this scenario is

4.111

For the F54​ Model, the present value in this scenario is

0.507

and the present value of annuity in this scenario is

4.1114.111​.)

B14 Model

F54 Model

Investment

$ 260,000

$230,000

Useful life​ (years)

6

6

Estimated annual net cash inflows for useful life

$90,000

$35,000

Residual value

$40,000

$9,000

Depreciation method

Straight−line

Straight−line

Required rate of return

12​%

12​%

What is the total present value of future cash inflows and residual value from the B14​ Model?
Answer #1
Year Present value factor (i=12%,n=6) 1/(1+0.12)6 =4.1111 Net cash inflow Present value of cash inflows =Present value factor *net cash inflow
1-6 Present value of annuity    4.1111 $90,000 4.1111*$90,000 = $3,69,999
6 Residual value/ salvage value 0.507   $40,000 0.507*40,000 = 20,280
Total present value of net cash inflows $3,90,279
Minus Investment ( $2,60000)

Net present value    $1,30,279

or

Total present value for future cash inflows of B14 model

Residual value = 0.507 * 40,000 =$20,280 Pv factor = 0.507

Depreciation of B14 model = $2,60,000-$40,000 / 6 =36,667